OpenStudy (anonymous):

simplify sqrt 8 (times) sqrt 32

OpenStudy (watchmath):

\(\sqrt{8}\sqrt{32}=\sqrt{8}\sqrt{8\cdot 4}=8\sqrt{4}=16\)